# A Survey of the Holroyd-Talbot Conjecture

### v1

@@ Abstract
A family of sets is intersecting if every pair of its members has an element in common. Such a family of sets is called a star if some element is in every set of the family. Given a graph $G$, let ${\mu}(G)$ denote the size of the smallest maximal independent set of $G$. In 2005, Holroyd and Talbot conjectured the following generalization of the Erdős-Ko-Rado Theorem: if $21\leq r\leq{\mu}(G)/$ 2 then there is an intersecting family of independent $r$-sets of maximum size that is a star. In this paper we present the history of this conjecture and survey the results that have supported it over the last 20 years.
@@ 1 Introduction
Since the publication of Theorem 1, in which the authors used shifting to achieve the result, a number of different methods have been used to prove it: bounding the size of the largest intersecting non-star family via shifting [38]; finding local bounds that extend to global bounds via the cycle method [51]; applying the Kruskal-Katona Theorem to shadows of complements [18, 50]; providing an upper bound on the Shannon capacity of the Kneser graph [56]; using Delsarte’s linear programming bound [66]; applying the Hoffman-Delsarte ratio bound to the independence number of the Kneser graph [36]; using the method of linearly independent polynomials [35]; and displaying an injection from any intersecting family into a star [31, 48].
For a graph $G$, we write ${\cal I}(G)$ for the family of all independent sets of $G$. Suppose that ${\cal F}\subseteq{\cal I}^{r}(G)$ is an intersecting subfamily of maximum size. Define $G$ to be $r$-EKR if some $v$ satisfies $|{\cal I}^{r}_{v}(G)|=|{\cal F}|$, and strictly $r$-EKR if every such ${\cal F}$ equals ${\cal I}_{v}^{r}(G)$ for some $v$. Such a $v$ is called an $r$-witness of $G$ (either strictly so or not). Let ${\alpha}(G)$ denote the independence number of $G$, and ${\mu}(G)$ denote the size of a smallest maximal independent set of $G$; i.e., the size of the smallest independent dominating set of $G$. Holroyd and Talbot [45] made the following conjecture to generalize the $t=1$ case of Theorem 1, which is the empty graph case of this conjecture.
@@
(Holroyd-Talbot, 2005) For any graph $G$, if $21\leq r\leq{\mu}(G)/$ 2 then $G$ is $r$-EKR, and strictly so if $21<r<{\mu}(G)/$ 2.
The necessity of the ${\mu}(G)/2$ condition in this conjecture stems from their example of two disjoint [^1] When we speak of disjoint graphs, we mean vertex-disjoint. copies of the complete bipartite graph $G=K_{3,3}=(A\times B)\cup(C\times D)$. Here we have ${\mu}(G)=6$ and a maximum star in ${\cal I}^{4}(G)$ has size 20, while the family consisting of every 4-set with at least two elements of $A$ is an intersecting non-star of size 24. Hence $G$ is not $4$-EKR. More generally, two disjoint copies of $K_{2k-1,2k-1}$ has ${\mu}/2=2k-1$ and is not $2k$-EKR. Maybe the simplest example is the $k$-pizza: the bipartite graph formed from the cycle $C_{2k}$ by adding the perfect matching between opposite pairs of vertices. The 5-pizza has ${\mu}=5$, and the family of all $\binom{5}{3}=10$ triples from one partite set is intersecting, while a 3-star only has size 9.
@@ 2 Precursors to the Holroyd-Talbot Conjecture
Deza and Frankl [21] considered the following scenario. An $\ell$-signed set is a pair $(A,f)$, where $A\ {\subseteq}\ [n]$ and $f:A{\rightarrow}[l]$, and its size equals $|A|$. For two such pairs $(A,f)$ and $(B,g)$ define $(A,f)\cap(B,g)=(C,h)$, where $C=\{x\in A\cap B\mid f(x)=g(x)\}$ and $h:C{\rightarrow}[\ell]$ such that $h(x)=f(x)$. Say that a family is $t$-intersecting if the size of the intersection of any two of its members is at least $t$. They proved the following, first stated by Meyer in [57] (stated in the complementary form of Berge/Livingston).
@@
If ${\cal F}\ {\subseteq}\ 2^{[n]}$ is intersecting then $|{\cal F}|\leq 2^{n-1}$.
Fact 5 holds because at most one of each pair $\{F,{\overline{F}}\}$ can be a member of ${\cal F}$, and is best possible because of $2^{[n]}_{x}$. Thus, if $A\in{\cal F}$ and ${\cal F}$ is an intersecting family of 2-signed sets, then there are at most $\binom{n}{r}$ $r$-sets $A$ in some pair of ${\cal F}$ when $r\leq n/2$ by Theorem 1. In addition, for each such $A$, Fact 5 implies that there are at most $2^{r-1}$ $f$ such that $(A,f)\in{\cal F}$, since the set of those $f$ is isomorphic to the set of characteristic vectors of an intersecting family of subsets of $A$. But Theorem 4 needs to be verified for the wider range $r\leq n$, as well as for $t>1$ and for $\ell>2$. Deza and Frankl prove this by using the shifting technique introduced in [24], which we define now.
For $1\leq i<j\leq n$, and for any family ${\cal F}\ {\subseteq}\ 2^{[n]}$, define the shifting function ${\sigma}_{j,i}:{\cal F}{\rightarrow}2^{[n]}$ by ${\sigma}_{j,i}(A)=A+i-j$ if $j\in A$ and $i\not\in A$, and $A+i-j\not\in{\cal F}$, and ${\sigma}_{j,i}(A)=A$ otherwise. We extend the shifting function to families of sets by ${\sigma}_{j,i}({\cal F})=\{{\sigma}_{j,i}(A)\mid A\in{\cal F}\}$. It is well known that $|{\sigma}_{j,i}({\cal F})|=|{\cal F}|$ and that ${\sigma}_{j,i}({\cal F})$ is $t$-intersecting whenever ${\cal F}$ is. Thus, when proving EKR-type results, one may always assume that an intersecting family ${\cal F}$ is shifted; that is, ${\sigma}_{j,i}({\cal F})={\cal F}$ for all $1\leq i<j\leq n$. (See [29] for the many other wonderful properties and implications of shifting.) The shifting technique as been applied to many variaties of intersection theorems, as well as to other types of problems in extremal set theory and other areas of combinatorics and graph theory. Its power and utility makes it a primary tool in tackling Conjecture 2.
The reason that Theorem 4 is a precursor to Conjecture 2 is that the family of all $\ell$-signed $r$-subsets of $[n]$ corresponds to ${\cal I}^{r}(G)$, where $G$ is the disjoint union of $n$ copies of the complete graph $K_{\ell}$, which we denote $nK_{\ell}$. With this correspondence, we may think of the $t=1$ case of Theorem 4 as being the second verification of Conjecture 2. In fact, this result goes farther than Holroyd-Talbot because it holds true for $r\leq n={\mu}(nK_{\ell})$. They also showed that $nK_{\ell}$ is strictly $r$-EKR unless $\ell=2$ and $r=n\geq 3$.
@@
    if $n\geq 2p+2s$ then ${\cal H}^{(p,s)}(n)$ is $(2p+s)$-EKR, and strictly so when $n>2p+2s$,
@@
Another scenario involves families of subsets of integers, in which no subset contains consecutive elements. Such subsets of $[n]$ correspond to independent sets in the path $P_{n}$. Disallowing the additional pair $\{1,n\}$ corresponds to independent sets in the cycle $C_{n}$; such subsets are known as separated sets. More generally, for a graph $G$, denote by $G^{(k)}$ the ${k^{\rm th}}$ power of $G$: the graph on $V(G)$ having an edge $xy$ precisely when the distance between $x$ and $y$ is positive and at most $k$. Then a $k$-separated set is any set in ${\cal I}(G^{(k)})$. Talbot [65] proved the following result, stated in terms of $k$-separated sets, which had been conjectured by Holroyd and Johnson in [43] (originally posed by Holroyd [42] in terms of sorties of King Arthur’s knights sitting far apart from one another at the Round Table).
@@
Talbot’s proof uses induction, a simplified shift operator, a very careful partition of ${\cal F}$ into $k+3$ parts, and seven pages of detailed analysis. When $k=1$ and $n=2r+2$, a 1-centered $r$-star in $C_{n}^{(k)}$ has size $|{\cal I}^{r-1}(P_{2r-1})|=\binom{r+1)}{r-1}=\binom{r+1}{2}$, and an example of a non-star intersecting family of that size is given by the following construction. Call a vertex of $V=V(C_{2r+2})$ even (odd) if its index is even (resp. odd). For any $1\leq i\leq r$, any set $S$ of $i$ consecutive even vertices, there is a unique set $T=T(S)$ of $r-i$ odd vertices such that $F_{S}=S\cup T\in{\cal I}^{r}(C_{2r+2})$. Let ${\cal S}$ denote the family of such $S$ for $i>(r+1)/2$; if $r$ is odd, also include half of the sets $S$ for $i=(r+1)/2$ by choosing one from each pair $\{S,{\overline{S}}\}$, where the complement is over even vertices. Then ${\cal S}$ is intersecting of size $\binom{r+1}{2}$; hence =${\cal F}=\{F_{S}\mid S\in{\cal S}\}$ is intersecting of that size.
@@
Theorem 10 follows from Theorem 8 by a straightforward lemma they prove: If $G$ is $r$-EKR with center $v$ then, for any set $S\ {\subseteq}\ N_{G}(v)$ of neighbors of of $v$, $G-S$ is also $r$-EKR with center $v$. The trick in applying the lemma is to remove, not all neighbors of $v$ but, rather, just those on the same side of $v$.
@@ 3 Further Support for the Holroyd-Talbot Conjecture
    If ${\alpha}(G)=2$ then $G$ is strictly $2$-EKR.
@@
    If ${\alpha}(G)\geq 3$ then $G$ is $2$-EKR if and only if ${\delta}\leq n-4$, and strictly so if and only if ${\delta}\leq n-5$, with any vertex of minimum degree as a maximum center.
@@
They achieved this result by (a) exhibiting conditions that produce an $r$-witness and (b) using those conditions to show that if $v$ is an $r$-witness of $G$ then $(v,x)$ is an $r$-witness of $G[H]$ for all $x\in V(K_{m})$.
@@
In [8], Borg had already generalized Conjecture 2 to include all hereditary families. In [9], he generalized it further, conjecturing the strengthening of Theorem 16 that it holds with $m(r,t)=(t+1)(r-t+1)$, adding also that the conclusion holds strictly unless ${\mu}({\cal F})=m(r,t)$ and $S=\{r\}$. He proved that this strengthening is true in the case that ${\cal F}$ is shifted.
Also in 2009, Borg and Holroyd [14] expanded upon Theorem 11 significantly, as follows.
@@
If $G$ is a chordal graph with at least one singleton then $G$ is $r$-EKR for all $r\leq{\alpha}(G)/2$.
@@
In [47] we also find the verification of Conjecture 2 for a connected sequence of complete graphs called chains. For a sequence of complete graphs $K_{a_{1}},\ldots,K_{a_{m}}$, we form the chain graph $G=G(a_{1},\ldots,a_{m})$ by allowing two consecutive $K_{a_{i}}$ and $K_{a_{i+1}}$ to share a single, distinct vertex, which requires of course that each $a_{i}\geq 2$. The chain $G$ is called special if $m\leq 2$ or if $m>2$, $a_{i}>a_{i-1}$ for each $2\leq i<m$, and $a_{m}\geq a_{m-1}$. They proved the following.
@@
If $G$ is a disjoint union of at most two special chains then $G$ is $r$-EKR for all $r$.
@@
If $G$ is a disjoint union of two cycles then $G$ is $r$-EKR for all $r$; furthermore, for any of its cycles that is even, all its vertices are $r$-witnesses.
@@
Let $G$ be a disjoint union of powers of cycles and a power of $H$, where $H$ is either a path or a cycle. Write $G=H^{k}\cup_{i=1}^{m}C_{n_{i}}^{k_{i}}$ and suppose that $k\leq min_{i}k_{i}$. Then $G$ is ${\alpha}(G)$-EKR for all $r$, and some $r$-witness has minimum degree.
@@
One of the simplest open cases for disjoint unions of graphs was the union of $k$ length-2 paths, denoted $kP_{3}$. Generalizing Theorem 4 (on $dK_{2}=kP_{2}$) in this graph manner, this case was partially resolved in 2020 [27] by also generalizing Bollobás and Leader’s use of the idea behind Katona’s cycle method.
@@
If $G$ is a graph on $n$ vertices with ${\Delta}(G)\leq d$ then $G$ is $r$-EKR for $r<(8n/27d)^{1/2}$.
@@
If $G$ is a graph on $n$ vertices with ${\overline{d}}(G)\leq c$, for any $c\geq e/36$ then $G$ is $r$-EKR for $r<(n/18c)^{1/3}$.
@@
A split vertex is a graph of degree at least 3. A spider is a tree with exactly one split vertex.
@@
If $G$ is a spider on $n$ vertices then $G$ is $r$-EKR for $r\leq(n\ln 2)^{1/2}-((\ln 2)/2)^{1/2}$.
@@
The key to these theorems is the method of diversity (originally called unbalance in [53]), which for a family ${\cal F}$ is defined to be ${\gamma}({\cal F})=|{\cal F}|-|{\cal F}_{z}|$ for any maximum center $z$. The origins of the concept are found in a result of [22] that states all but at most $O(\binom{n-2}{k-2})$ members of any intersecting ${\cal F}\ {\subseteq}\binom{[n]}{r}$ contain a common element. The key lemma in [32] is a result from [30] that every intersecting ${\cal F}\ {\subseteq}\binom{[n]}{r}$, with $n>72r$, has ${\gamma}({\cal F})\leq\binom{n-3}{r-2}$. How this is used to prove Theorem 25, for example, is to let $z$ be a maximum center of ${\cal F}$ and choose some $E\in{\cal F}-{\cal F}_{z}$. Because $G$ has bounded degree $d$, one can compute a simple lower bound $f(n,r,d)$ on the number of sets $F\in{\cal I}_{z}^{r}(G)$ that are disjoint from $E$ and so can’t be in ${\cal F}$. Then $|{\cal F}|=|{\cal F}_{z}|+{\gamma}({\cal F})\leq|{\cal I}_{z}^{r}(G)|-f(n,r,d)+\binom{n-3}{r-2}$, which is at most $|{\cal I}_{z}^{r}(G)|$ when $f(n,r,d)\geq\binom{n-3}{r-2}$. Hence the bound on $r$ in the theorem come from making the last inequality true. The specific properties of the graphs in the other theorems are used to modify their bounds on $r$ similarly.
Another theorem on trees was given by Feghali, et al. [28]. A spider is $\ell$-uniform if every leaf has distance $\ell$ from the split vertex; it has also been called a $-\ell$ claw in the literature. For example, the 1-uniform spider with $k$ leaves is $K_{1,k}$, which is $r$-EKR for all $r\leq k/2$ by Theorem 1.
@@
Finally, Holroyd and Talbot [45] opined that, if a counterexample $G$ to Conjecture 2 exists, then one exists with ${\mu}(G)={\alpha}(G)$. This motivates studying such graphs, which are equivalent to the well-known class of well-covered graphs [^3] A graph is well-covered if all its minimal vertex covers have the same size. A vertex cover is a set of vertices that touch all edges.. Examples of such graphs already studied above include disjoint unions of complete graphs, balanced complete bipartite graphs, and graphs of independence number 2 with no dominating vertex (e.g., complements of cycles). A pendant vertex has degree 1. Another construction of a well-covered graph is a fully pendant graph $G^{*}$, which is formed from a graph $G$ by adding to each of its vertices a pendant neighbor. De Silva, et al. [19] recently proved the following.

### v3

@@ Abstract
A family of sets is intersecting if every pair of its members has an element in common. Such a family of sets is called a star if some element is in every set of the family. Given a graph $G$, let ${\mu}(G)$ denote the size of the smallest maximal independent set of $G$. In 2005, Holroyd and Talbot conjectured the following generalization of the Erdős-Ko-Rado Theorem: for $21\leq r\leq{\mu}(G)/$ 2, there is a maximum size intersecting family of independent $r$-sets that is a star. In this paper we present the history of this conjecture and survey the results that have supported it over the last 20 years.
@@ 1 Introduction
Since the publication of Theorem 1, in which the authors used shifting to achieve the result, a number of different methods have been used to prove it: bounding the size of the largest intersecting non-star family via shifting [38]; finding local bounds that extend to global bounds via the cycle method [52]; applying the Kruskal-Katona Theorem to shadows of complements [18, 51]; providing an upper bound on the Shannon capacity of the Kneser graph [57]; using Delsarte’s linear programming bound [68]; applying the Hoffman-Delsarte ratio bound to the independence number of the Kneser graph [36]; using the method of linearly independent polynomials [35]; displaying an injection from any intersecting family into a star [31, 48]; and using the Borel Fixed Point Theorem from algebraic group theory [70].
For a graph $G$, we write ${\cal I}(G)$ for the family of all independent sets of $G$, and use ${\cal I}^{r}(G)$ in place of ${\cal I}(G)^{r}$ and ${\cal I}_{v}(G)$ in place of ${\cal I}(G)_{v}$, etc. Suppose that ${\cal F}\subseteq{\cal I}^{r}(G)$ is an intersecting subfamily of maximum size. Define $G$ to be $r$-EKR if some $v\in V(G)$ satisfies $|{\cal I}^{r}_{v}(G)|=|{\cal F}|$, and strictly $r$-EKR if every such ${\cal F}$ equals ${\cal I}_{v}^{r}(G)$ for some $v$. Such a $v$ is called an $r$-witness of $G$ (either strictly so or not). Let ${\alpha}(G)$ denote the independence number of $G$, and ${\mu}(G)$ denote the size of a smallest maximal independent set of $G$; i.e., the size of the smallest independent dominating set of $G$. Holroyd and Talbot [45] made the following conjecture to generalize the $t=1$ case of Theorem 1, which is the empty graph case of this conjecture.
@@
(Holroyd-Talbot, 2005) For any graph $G$, if $21\leq r\leq{\mu}(G)/$ 2, then $G$ is $r$-EKR, and strictly so if $21<r<{\mu}(G)/$ 2.
The necessity of the ${\mu}(G)/2$ condition in this conjecture stems from their example of two disjoint [^1] When we speak of disjoint graphs, we mean vertex-disjoint. copies of the complete bipartite graph $K_{3,3}$: $G=2K_{3,3}$, with $E(G)=(A\times B)\cup(C\times D)$. Here we have ${\mu}(G)=6$ and a maximum star in ${\cal I}^{4}(G)$ has size 20, while the family consisting of every independent 4-set with at least two elements of $A$ is an intersecting non-star of size 24. Hence $G$ is not $4$-EKR. More generally, two disjoint copies of $K_{2k-1,2k-1}$ has ${\mu}/2=2k-1$ and is not $2k$-EKR.
@@ 2 Precursors to the Holroyd-Talbot Conjecture
Deza and Frankl [21] considered the following scenario. An $\ell$-signed set is a pair $(A,f)$, where $A\ {\subseteq}\ [n]$ and $f:A{\rightarrow}[\ell]$, and its size equals $|A|$. For two such pairs $(A,f)$ and $(B,g)$ define $(A,f)\cap(B,g)=(C,h)$, where $C=\{x\in A\cap B\mid f(x)=g(x)\}$ and $h:C{\rightarrow}[\ell]$ such that $h(x)=f(x)$. Say that a family of $\ell$-signed sets is $t$-intersecting if the size of the intersection of any two of its members is at least $t$. They proved the following, first stated by Meyer in [58] (stated in the complementary form of Berge/Livingston).
@@
If ${\cal F}\ {\subseteq}\ 2^{[n]}$ is intersecting, then $|{\cal F}|\leq 2^{n-1}$.
Fact 5 holds because at most one of each pair $\{F,{\overline{F}}\}$ can be a member of ${\cal F}$, and is best possible because of $2^{[n]}_{x}$. Thus, if ${\cal F}$ is an intersecting family of 2-signed sets, then there are at most $\binom{n-1}{r-1}$ $r$-sets $A$ in some pair of ${\cal F}$ when $r\leq n/2$ by Theorem 1. In addition, for each such $A$, Fact 5 implies that there are at most $2^{r-1}$ functions $f$ such that $(A,f)\in{\cal F}$, since the set of those $f$ is isomorphic to the set of characteristic vectors of an intersecting family of subsets of $A$. That is, for fixed $A\in\binom{[n]}{r}$, the set of pairs $(A,f)$, where $f:A{\rightarrow}[2]$ is isomorphic to $2^{A}$ by the correspondence $B(A,f)=\{a\in A\mid f(a)=2\}$; clearly $(A,f)\cap(A^{\prime},f^{\prime})\not={\emptyset}$ if and only if $B(A,f)\cap B(A^{\prime},f^{\prime})\not={\emptyset}$. But Theorem 4 needs to be verified for the wider range $r\leq n$, as well as for $t>1$ and for $\ell>2$. Deza and Frankl [21] prove this by using the shifting technique introduced in [24], which we define now.
For $1\leq i<j\leq n$, and for any family ${\cal F}\ {\subseteq}\ 2^{[n]}$, define the shifting function ${\sigma}_{j,i}:{\cal F}{\rightarrow}2^{[n]}$ by
${\sigma}_{j,i}(A)=\left\{\begin{array}[]{ll}A+i-j&\mbox{if}\ j\in A,\ i\not\in A,\ \mbox{and}\ A+i-j\not\in{\cal F},\ \mbox{and}\\ {\sigma}_{j,i}(A)=A&\mbox{otherwise.}\end{array}\right.$
We extend the shifting function to families of sets by ${\sigma}_{j,i}({\cal F})=\{{\sigma}_{j,i}(A)\mid A\in{\cal F}\}$. It is well known that $|{\sigma}_{j,i}({\cal F})|=|{\cal F}|$ and that ${\sigma}_{j,i}({\cal F})$ is $t$-intersecting whenever ${\cal F}$ is. Thus, when proving EKR-type results, one may always assume that an intersecting family ${\cal F}$ is shifted; that is, ${\sigma}_{j,i}({\cal F})={\cal F}$ for all $1\leq i<j\leq n$. (See [29] for the many other wonderful properties and implications of shifting.) The shifting technique as been applied to many varieties of intersection theorems, as well as to other types of problems in extremal set theory and other areas of combinatorics and graph theory. Its power and utility makes it a primary tool in tackling Conjecture 2.
The reason that Theorem 4 is a precursor to Conjecture 2 is that the family of all $\ell$-signed $r$-subsets of $[n]$ corresponds to ${\cal I}^{r}(G)$, where $G$ is the disjoint union of $n$ copies of the complete graph $K_{\ell}$, which we denote $nK_{\ell}$. With this correspondence, we may think of the $t=1$ case of Theorem 4 as being the second verification of Conjecture 2. In fact, this result goes farther than Holroyd-Talbot because it holds true for $r\leq n={\mu}(nK_{\ell})$. Deza and Frankl [21] also showed that $nK_{\ell}$ is strictly $r$-EKR unless $\ell=2$ and $r=n\geq 3$.
@@
    if $n\geq 2p+2s$, then ${\cal H}^{(p,s)}(n)$ is $(2p+s)$-EKR, and strictly so when $n>2p+2s$,
@@
Another scenario involves families of subsets of integers, in which no subset contains consecutive elements. Such subsets of $[n]$ correspond to independent sets in the path $P_{n}$. Disallowing the additional pair $\{1,n\}$ corresponds to independent sets in the cycle $C_{n}$; such subsets are known as separated sets. More generally, for a graph $G$, denote by $G^{(k)}$ the ${k^{\rm th}}$ power of $G$: the graph on $V(G)$ having an edge $xy$ precisely when the distance in $G$ between $x$ and $y$ is positive and at most $k$. Then a $k$-separated set is any set in ${\cal I}(C_{n}^{(k)})$. Talbot [67] proved the following result, stated in terms of $k$-separated sets, which had been conjectured by Holroyd and Johnson in [43] (originally posed by Holroyd [42] in terms of sorties of King Arthur’s knights sitting far apart from one another at the Round Table).
@@
Talbot’s proof uses induction, a simplified shift operator, a very careful partition of ${\cal F}$ into $k+3$ parts, and seven pages of detailed analysis. When $k=1$ and $n=2r+2$, a 1-centered $r$-star in $C_{n}^{(k)}$ has size $|{\cal I}^{r-1}(P_{2r-1})|=\binom{r+1}{r-1}=\binom{r+1}{2}$, and an example of a non-star intersecting family of that size is given by the following construction. Call a vertex of $V=V(C_{2r+2})$ even (odd) if its index is even (resp. odd). For any $1\leq i\leq r$, any set $S$ of $i$ consecutive even vertices, there is a unique set $T=T(S)$ of $r-i$ odd vertices such that $F_{S}=S\cup T\in{\cal I}^{r}(C_{2r+2})$. Let ${\cal S}$ denote the family of such $S$ for $i>(r+1)/2$; if $r$ is odd, also include half of the sets $S$ for $i=(r+1)/2$ by choosing one from each pair $\{S,{\overline{S}}\}$, where the complement is over even vertices. Then ${\cal S}$ is intersecting of size $\binom{r+1}{2}$; hence =${\cal F}=\{F_{S}\mid S\in{\cal S}\}$ is intersecting of that size.
@@
Theorem 10 follows from Theorem 8 by a straightforward lemma that they prove: If $G$ is $r$-EKR with center $v$ then, for any set $S\ {\subseteq}\ N_{G}(v)$ of neighbors of of $v$, $G-S$ is also $r$-EKR with center $v$. The trick in applying the lemma is to remove, not all neighbors of $v$ but, rather, just those on the clockwise side of $v$.
@@ 3 Further Support for the Holroyd-Talbot Conjecture
    If ${\alpha}(G)=2$, then $G$ is strictly $2$-EKR.
@@
    If ${\alpha}(G)\geq 3$, then $G$ is $2$-EKR if and only if ${\delta}\leq n-4$, and strictly so if and only if ${\delta}\leq n-5$, with any vertex of minimum degree as a maximum center.
@@
They achieved this result by (a) exhibiting conditions that produce an $r$-witness and (b) using those conditions to show that if $v$ is an $r$-witness of $G$ then $(v,x)$ is an $r$-witness of $G[K_{m}]$ for all $x\in V(K_{m})$.
@@
In [8], Borg had already generalized Conjecture 2 to include all hereditary families. In [9], he generalized it further, conjecturing the strengthening of Theorem 16 that it holds with $m(r,t)=(t+1)(r-t+1)$, adding also that the conclusion holds strictly unless ${\mu}({\cal F})=m(r,t)$ and $S=\{r\}$. He proved that this strengthening is true for shifted families ${\cal F}$.
Also in 2009, Borg and Holroyd [14] expanded upon Theorem 11 significantly, as follows. They call a graph a modified power of a cycle if it can be obtained from some $C_{n}^{(k)}$ by adding a set of consecutive edges of distance $k+1$ in $C_{n}$.
@@
If $G$ is a chordal graph with at least one singleton, then $G$ is $r$-EKR for all $r\leq{\alpha}(G)/2$.
@@
Soon after, Woodroofe [69] proved a more general result for hereditary families via (exterior) algebraic shifting, as opposed to the combinatorial shifting discussed above. Algebraic shifting was introduced by Kalai [50] and takes advantage of the geometry that arises from viewing a hereditary family as an abstract simplicial complex. For example, the concepts of a simplicial elimination (or dismantling) ordering for chordal (or dismantlable) graphs generalize to shellability and the even more general sequentially Cohen-Macauley property for simplicial complexes, which we will not define here. A graph $G$ is called sequentially Cohen-Macaulay if ${\cal I}(G)$ is sequentially Cohen-Macaulay. Among the graphs in this class are chordal graphs, graphs with no induced cycles of length other than 3 or 5, incomparability graphs of shellable posets, and others (see [69]). Woodroofe proved the following.
@@
If $G$ is a sequentially Cohen-Macauley graph with an isolated vertex, then $G$ is $r$-EKR for all $r\leq{\mu}(G)$.
Additionally, Woodroofe uses these tools to prove the most general version to date along the lines of Theorem 17.
Theorem 20.
Let $G$ be the disjoint union of $k$ graphs, including a singleton. Then $G$ is $r$-EKR for all $r\leq k/2$.
Regarding graphs without singletons, in [47] we also find the verification of Conjecture 2 for a connected sequence of complete graphs called chains. For a sequence of complete graphs $K_{a_{1}},\ldots,K_{a_{m}}$, we form the chain graph $G=G(a_{1},\ldots,a_{m})$ by allowing two consecutive $K_{a_{i}}$ and $K_{a_{i+1}}$ to share a single, distinct vertex, which requires of course that each $a_{i}\geq 2$. The chain $G$ is called special if $m\leq 2$ or if $m>2$, $a_{i}>a_{i-1}$ for each $2\leq i<m$, and $a_{m}\geq a_{m-1}$. They proved the following.
Theorem 21.
If $G$ is a disjoint union of at most two special chains, then $G$ is $r$-EKR for all $r$.
@@
If $G$ is a disjoint union of two cycles then $G$ is $r$-EKR for all $r$; furthermore, all of the vertices of any of its even cycles are $r$-witnesses.
@@
Let $G$ be a disjoint union of powers of cycles and a power of $H$, where $H$ is either a path or a cycle. Write $G=H^{k}\cup\left(\cup_{i=1}^{m}C_{n_{i}}^{k_{i}}\right)$ and suppose that $k\leq min_{i}k_{i}$. Then $G$ is ${\alpha}(G)$-EKR for all $r$, and some $r$-witness has minimum degree.
@@
One of the simplest open cases for disjoint unions of graphs was the union of $k$ length-2 paths, denoted $kP_{3}$. Generalizing Theorem 4 (on $kK_{2}=kP_{2}$) in this manner, this case was partially resolved in 2020 [27] by also generalizing Bollobás and Leader’s use of the idea behind Katona’s cycle method.
@@
If $G$ is a graph on $n$ vertices with ${\Delta}(G)\leq d$, then $G$ is $r$-EKR for $r<(8n/27d)^{1/2}$.
@@
If $G$ is a graph on $n$ vertices with ${\overline{d}}(G)\leq c$, for any $c\geq e/36$, then $G$ is $r$-EKR for $r<(n/18c)^{1/3}$.
@@
A split vertex is a vertex of degree at least 3. A spider is a tree with exactly one split vertex.
@@
If $G$ is a spider on $n$ vertices, then $G$ is $r$-EKR for $r\leq(n\ln 2)^{1/2}-((\ln 2)/2)^{1/2}$.
@@
The key to these theorems is the method of diversity (originally called unbalance in [54]), which for a family ${\cal F}$ is defined to be ${\gamma}({\cal F})=|{\cal F}|-|{\cal F}_{z}|$ for any maximum center $z$. The origins of the concept are found in a result of [22] that states all but at most $O(\binom{n-2}{k-2})$ members of any intersecting ${\cal F}\ {\subseteq}\binom{[n]}{r}$ contain a common element. The key lemma in [32] is a result from [30] that every intersecting ${\cal F}\ {\subseteq}\binom{[n]}{r}$, with $n>72r$, has ${\gamma}({\cal F})\leq\binom{n-3}{r-2}$. How this is used to prove Theorem 27, for example, is to let $z$ be a maximum center of ${\cal F}$ and choose some $E\in{\cal F}-{\cal F}_{z}$. Because $G$ has bounded degree $d$, one can compute a simple lower bound $f(n,r,d)$ on the number of sets $F\in{\cal I}_{z}^{r}(G)$ that are disjoint from $E$ and so can’t be in ${\cal F}$. Then $|{\cal F}|=|{\cal F}_{z}|+{\gamma}({\cal F})\leq|{\cal I}_{z}^{r}(G)|-f(n,r,d)+\binom{n-3}{r-2}$, which is at most $|{\cal I}_{z}^{r}(G)|$ when $f(n,r,d)\geq\binom{n-3}{r-2}$. Hence the bound on $r$ in the theorem comes from making the last inequality true. The specific properties of the graphs in the other theorems are used to modify their bounds on $r$ similarly.
Another theorem on trees was given by Feghali, Hurlbert, and Kamat [28]. A spider is $\ell$-uniform if every leaf has distance $\ell$ from the split vertex; it has also been called a $-\ell$ claw in the literature. For example, the 1-uniform spider with $k$ leaves is $K_{1,k}$, which is $r$-EKR for all $r\leq k/2$ by Theorem 1.
@@
Finally, Holroyd and Talbot [45] opined that, if a counterexample $G$ to Conjecture 2 exists, then one exists with ${\mu}(G)={\alpha}(G)$. This motivates studying such graphs, which are equivalent to the well-known class of well-covered graphs: a vertex cover is a set of vertices that touch all edges, and a graph is well-covered if all its minimal vertex covers have the same size. Examples of such graphs already studied above include disjoint unions of complete graphs, balanced complete bipartite graphs, and graphs of independence number 2 with no dominating vertex (e.g., complements of cycles). A pendant vertex has degree 1. Another construction of a well-covered graph is a fully pendant graph $G^{*}$, which is formed from a graph $G$ by adding to each of its vertices a pendant neighbor. De Silva, Dionne, Dunkelberg, and Harris [19] recently proved the following.